This page left intentionally

blank

No, forget that This program was written to help you do pagination tricks that are difficult to do with TEX under normal circumstances. This program can be used for tasks as simple as reversing the order of pages in a dvi file or selecting the even or odd pages. It can be used to print little cards like this, or brochures, by combining multiple pages onto one page in some regular way. You can print folded booklets and many other things. Enjoy!

3